OpenStudy (anonymous):

Most athletic shoes have coefficients of static friction of 0.80 on a tile surface. What is the minimum time in which a person at rest could move 3 meters on this surface if they start from rest?

OpenStudy (anonymous):

wat is acc(deceleration) caused by frictional force?

OpenStudy (anonymous):

Fapp-f=ma?

OpenStudy (anonymous):

dont consider the force that u give to sustain its motion wat is the minimum time required to move 3m that is time taken while the decelerating force (acceleration that is) F=ma=0.8*mg so a=?

OpenStudy (anonymous):

So a=7.84m/s2? What about the 3 meters?

OpenStudy (anonymous):

consider the frictional force to be uniform and use the eqns for uniform acceleratiun u know them?

OpenStudy (anonymous):

uhh do you mean: fsmax=\[\mu s * \eta\] and fk=\[\mu k * \eta\]

OpenStudy (anonymous):

u r right F-Ff=ma a-a0=A(net) where F is the force u give to more it Ff is frictional force=mg*mui a0=g*mui find net acc Anet u know eqns 4 uniform acc?